OpenStudy (anonymous):

What is the domain and range of the quadratic equation y = –3(x – 10)2 + 5?

OpenStudy (anonymous):

domain of any polynomial function is all real numbers

OpenStudy (anonymous):

the range will be from negative infinity to the y-value of the vertex which is y=5; range is\[(-\infty,5)\]

OpenStudy (anonymous):

domain is (-infinity,infinity)

OpenStudy (anonymous):

domain is \[\mathbb{R}\]

OpenStudy (anonymous):

is it less than or greater than 5

jimthompson5910 (jim_thompson5910):

You can include the value y = 5 in the range, so the range is actually \[\Large (-\infty,5]\]

OpenStudy (anonymous):

^

OpenStudy (anonymous):

i am corrected

OpenStudy (anonymous):

sloppy on my part